OpenStudy (anonymous):

6 3/8 divided by 2 3/7??any help??

jimthompson5910 (jim_thompson5910):

\[\large \frac{6\frac{3}{8}}{2\frac{3}{7}}\] \[\large \frac{\frac{8*6+3}{8}}{\frac{7*2+3}{7}}\] \[\large \frac{\frac{51}{8}}{\frac{17}{7}}\] \[\large \frac{51}{8}\frac{7}{17}\] \[\large \frac{51*7}{8*17}\] \[\large \frac{357}{136}\] \[\large \frac{21}{8}\] \[\large \frac{16+5}{8}\] \[\large \frac{16}{8}+\frac{5}{8}\] \[\large 2+\frac{5}{8}\] \[\large 2\frac{5}{8}\] So \[\large \frac{6\frac{3}{8}}{2\frac{3}{7}}=2\frac{5}{8}\]

OpenStudy (anonymous):

much thanx